a swimming pool has an area of 40 square meters and a perimeter of 26 meters. What is the length and width of the swimming pool? Is it possible for another pool to have the same area but a larger perimeter ?
I think the length would be 10meters and 3 meters wide ? not sure how to do the rest

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

a swimming pool has an area of 40 square meters and a perimeter of 26 meters. What is the length and width of the swimming pool?
10 by 3 = 30 sq meters, not 40.
It's 5 by 8
-----------------
Is it possible for another pool to have the same area but a larger perimeter ?
Yes, it is.
It could be 0.1 m by 400 meter --> P = 800.2 meters
--------------------
L*W = 40
2L + 2W = 26
--> L = 8, W = 5 or vice versa


Answer by stanbon(75887)

a swimming pool has an area of 40 square meters and a perimeter of 26 meters. What is the length and width of the swimming pool?
Perimeter = 2(L+W)
26 = 2(L+W)
13 = L+W
L = 13-W
----
Area = L*W
40 = (13-W)W
40 = 13W-W^2
W^2 -13W+40 = 0
W = [13 +- sqrt(169-4*1*40)]/2
W = [13 +- sqrt(9)}/2
W = [13+-3]/2
Positive solution:
W = 8 meters
----
So L = 5 meters
====================
Is it possible for another pool to have the same area but a larger perimeter ?
Yes
--------------
Note A pool that is 40 meters by 1 meter has a perimeter
of (40+1+40+1) = 82 meters. This pool still has an area
of 40 sq. meters.
===================================
Cheers,
Stan H.
=====================
==========================================